# HACTS: a Human-As-Copilot Teleoperation System for Robot Learning

### v1

@@ I Introduction
However, current teleoperation systems are limited by their hardware, such as VR [9], exoskeletons [5], and motion-capture technologies [4], which are designed for unilateral control. In these systems, the human operator can send commands to the robot but receives no real-time feedback, a limitation when the robot is performing autonomous tasks and requires human intervention. Without feedback, the robot may make unintended, sudden movements, disrupting task continuity and performance. This is akin to a driver of an autonomous vehicle attempting to intervene while the vehicle is in motion—without a mechanism to sync the vehicle’s wheel movement with the driver’s input, the intervention could lead to a loss of control [26]. Recent work [10, 11, 19] has proposed bilateral control systems that integrate tactile and force feedback into teleoperation hardware. However, these systems require the follower device to generate precise force or tactile sensing information—capabilities that are not widely available on most robotic platforms. Thus, there is a clear need for a teleoperation system offering bilateral joint synchronization between the teleoperation system and the robot, ensuring smooth transitions and more reliable interventions, as well as being compatible with a broad range of robots.
To address this limitation, in this paper, we propose HACTS (Human-As-Copilot Teleoperation System), a low-cost but effective solution that enables bilateral, real-time joint synchronization between the robot arm and the teleoperation hardware. The HACTS system is implemented using only 3D-printed components and low-cost, off-the-shelf motors, making it both affordable and scalable. This simple yet effective hardware setup allows for continuous joint feedback from the robot to the teleoperation hardware, enabling more intuitive and controlled intervention when required. HACTS offers a robust, seamless connection between the operator and the robot arm, making it ideal for human-in-the-loop applications.
@@ II-A Teleoperation for robot manipulation learning
Despite recent advances in robot manipulation driven by AI, fully autonomous solutions remain far from achieving socially and physically competent robot behaviors. This is why teleoperated robots, serving as physical avatars for human workers on-site, present a more practical solution [4, 20]. Furthermore, with the rise of learning-based, data-driven approaches in robotic control, teleoperation has become an essential system for gathering manipulation data [23, 9]. Traditional teleoperation systems, such as VR [27], joystick controller [7], Phone-based [16], and 3D Spacemouse [13], can only control the end-effector pose of the robot arm that heavily relies on IK/FK and motion planning, which limits the expression of human operators. More recent approaches have highlighted the potential of low-cost, kinematically equivalent controllers, such as ALOHA [28] and Gello [25], which enable more dexterous and precise control, providing valuable data for robot learning. However, these systems have primarily focused on unilateral control, restricting the full potential of the hardware. Several works have attempted to develop bilateral control systems for robotic teleoperation [19, 11, 10], focusing on providing force or tactile feedback to human operators. However, these complex bilateral systems suffer from carefully tactile and force algorithms design, and require expensive follower hardware that can deliver accurate force or tactile information—features that are not accessible for most robotic platforms. In this work, we propose a similar low-cost, kinematically equivalent device that incorporates bilateral control, enabling humans to act as copilots and intervene when needed. We demonstrate that even a simple bilateral position synchronization, which is compatible with almost all robotic systems, can significantly enhance control precision and data collection capabilities. Our approach highlights the potential of more accessible, cost-effective teleoperation systems for improving robot learning and human-robot interaction.
@@ II-B Learning from human demonstration and correction
Learning from Demonstration (LfD) is a key paradigm in robot learning, enabling robots to acquire new skills by learning from human demonstrations rather than through explicit programming. Among the notable advancements in LfD are Vision-Language-Action (VLA) models [14, 2, 22, 21]. These approaches facilitate learning multimodal human demonstrations, increasingly leveraging larger model architectures with more parameters. As a result, robots can benefit from richer contextual information, which improves task performance and aligns with the scaling laws observed in large language models. However, the process of collecting human demonstrations can be expensive and time-consuming, with previous work primarily focusing on success trajectories [23, 9]. In contrast, our HACTS, with its simple bilateral control, allows for human intervention and generates more diverse types of data. Recent research has underscored the importance of failure data in training [24], and we believe our system can significantly contribute to advancing this line of work by supporting the collection of such critical data for more robust learning.
The other group of methods that can benefit from HACTS is Human-In-The-Loop (HITL) learning, which refines robot behavior through interactive feedback and enables policy adjustments based on human corrections. HITL methods can be categorized into passive human intervention [18, 12], where humans assist periodically or as determined by functions, and active human intervention [8, 17, 13], where humans decide when to intervene. Recently, approaches such as HACO [17] and HIL-SERL [13] leverage active human assistance to enhance RL exploration and ensure safety during training, improving sample efficiency and exhibiting robust behaviors compared to classical RL and imitation methods like HG-Dagger [8] and IWR [15]. However, these approaches require agile hardware platforms for flexible, timely intervention. In contrast to previous teleoperation systems, our HACTS offers more dexterous, precise control during human intervention, which we believe will inspire further innovation in HITL RL, especially for real-world online robot learning.
@@ III-A Overview
We illustrate the key advancements of HACTS in Figure 2. HACTS operates as a leader-follower system with bilateral position control, enabling both leader-to-follower and follower-to-leader synchronization. In leader-to-follower synchronization, HACTS functions similarly to systems like ALOHA [28] and Gello [25], facilitating the rapid and precise collection of offline demonstration data. Conversely, in follower-to-leader synchronization, where model inference drives the robot arm, HACTS enables the human operator to act as a copilot, seamlessly intervening when necessary to provide online action-correction data. Below, we outline key hardware and software design principles that guided the development of HACTS.
@@ III-B Hardware Design
Low-Cost. One of our primary goals is to demonstrate that a capable teleoperation system can be built at an affordable price, making it accessible for large-scale deployment. To achieve this, we opted for servo motors rather than more expensive robotic modules, which are commonly used in more complex systems [4]. The use of servo motors allows us to implement both passive position reading and active position control at a significantly lower cost. Previous works, including ALOHA [28], Gello [25], and FACTR [11], have shown the impressive performance and stability of low-cost DYNAMIXEL motors, which further supports our design choice.
@@
Easy-to-Use. Building upon prior work that demonstrated the effectiveness of miniature [25], kinematically equivalent devices for teleoperation, we aimed to create hardware that is intuitive and user-friendly, even for non-experts. This simplicity is key to ensuring that the system can be scaled for broader use, i.e., quickly adapt to a new robot. The kinematically equivalent structure is derived from the Denavit–Hartenberg (DH) parameters, with the lengths scaled by a factor to maintain the system’s usability while optimizing for lightweight design.
Note that these components can easily be extended to other robot arms since we only need to synchronize position information, which is generally available for most robotic arms.
@@ IV Experiments
We evaluate the efficacy of HACTS by addressing four key research questions:
Q1: Can human intervention data collected via the HACTS system serve as a viable substitute for an equivalent volume of direct teleoperation data?
Q2: Can the action-correction data in HACTS further improve the model’s success rate under in-distribution (ID) scenarios?
Q3: Can the action-correction data in HACTS enhance the model’s generalization capability for out-of-distribution (OOD) scenarios?
Q4: Can HACTS further support human-in-the-loop online reinforcement learning?
To investigate these questions, we developed real-world robotic arm environments, designed a range of manipulation tasks, and established diverse experimental setups incorporating two primary learning paradigms: imitation learning (IL) and reinforcement learning (RL).
@@ IV-A Hardware Setup
As shown in Figure 3, we built HACTS on two workstations, each equipped with one wrist-mounted Orbbec Gemini 336 camera, and one or multiple Orbbec Gemini 336L cameras. As illustrated in Figure 3, these cameras are positioned at different angles to capture a comprehensive set of visual inputs from varying perspectives. All components—including the cameras, the UR5 robot with Robotiq 2f-85 gripper, the HACTS system, and the policy model—are integrated into a single RTX 4090 workstation, ensuring robust computational performance for real-time data processing. To demonstrate the feasibility and effectiveness of HACTS, we implemented the hardware using the UR5 kinematic model, and reused parts of the structure design in [25] but modified them to fit the XL430 motors. The whole cost for HACTS hardware is lower than \$$\$300$, and we listed all components in Table I.
@@ IV-B Experimental Setup on Imitation Learning
Task Design. To evaluate the use of the HACTS within the framework of imitation learning, we designed three real-world robotic tasks. These tasks required the robotic arm to learn and execute complex manipulation actions, such as grasping, placing, pressing, flipping, and precisely identifying the location of objects. The objects used in these tasks are shown in Figure 4. The tasks are as follows: OpenBox: Accurately locate a switch of a box and press it to open the box. SteamBun: Grasp a steamed bun from a plate and place it into a steamer. UprightMug: Flip an overturned mug upright, then place it back in a designated central position.
Algorithm Instances. We focus on popular visuomotor policy learning techniques, where the robot learns to perform tasks based on visual observations and proprioception. The input consists of the RGB image data and the robot proprioceptive states, while the output is the end-to-end robotic control signals including 6-D joint positions and 1-D gripper value. Imitation learning methods allow the robot to learn directly from human demonstrations, mimicking the demonstrated actions to perform similar tasks autonomously. We employed two state-of-the-art imitation learning methods for our experiments:
  • Action Chunking Transformer (ACT) [28]: ACT predicts a sequence of actions and leverages temporal ensemble technology to generate smooth and precise actions.
  • Diffusion Policy (DP) [3]: DP uses a diffusion model, which is expressive and can model multi-modal action patterns, to improve the success rates.
Evaluation Protocol. For each task, we initially collected 100 expert trajectories through teleoperation. From these, we randomly chose 50 expert trajectories to train the preliminary ACT and DP models, referred to as “pre-ACT///DP”. Additionally, we trained the ACT and DP models using the full set of 100 expert trajectories as the baseline models, referred to as “full-ACT///DP”. Further, to answer the Q1, Q2 and Q3, we designed four different experimental settings. For each various experimental setting, we used HACTS to collect 50 human intervention trajectories (denoted as HACTS trajectories) with using the “pre-ACT///DP” to execute tasks. Then we mixed these 50 HACTS trajectories with the initial 50 random expert trajectories to train the refined ACT and DP models, denoted as “HACTS-ACT///DP”. Four experimental settings are as follows:
@@ IV-C Experimental Results on Imitation Learning
Table II presents the results of ACT on the three experimental settings including EADC, FCID, and ODSS for all three tasks, evaluated in terms of success rate. Table III presents the results of the DP algorithm on the setting ODDS.
Answer for Question 1. We focused on the comparison between HACTS-ACT and full-ACT in the EADC setting in Table II. HACTS-ACT achieved higher success rates than full-ACT in all three tasks. This indicates that the HACTS data can effectively replace expert teleoperation data without any performance drop. This aligns with our expectations, as both datasets are fundamentally derived from expert data collected through teleoperation systems.
Answer for Question 2. We examined whether the performance of HACTS-ACT improves in the FCID setting. The results presented in Table II reveal that HACTS-ACT not only significantly enhanced the success rates of pre-ACT across all three tasks, but also consistently outperformed full-ACT. For instance, in the UprightMug task, HACTS-ACT achieved a remarkable success rate of 70%, compared to just 10% for pre-ACT and 40% for full-ACT. This improvement can be attributed to the ability of HACTS-ACT to specifically gather corrective data targeted at failed cases, whereas full-ACT merely increases the volume of data.
Answer for Question 3. We analyzed the performance of HACTS-ACT under the ODSS setting in Table II and the ODDS setting in Table III. Notably, we found that the success rates of pre-ACT, pre-DP, full-ACT, and full-DP are all 0% in these scenarios, indicating that they are unable to handle situations that are unseen in the training dataset. However, upon incorporating HACTS-generated action-correction data, both HACTS-ACT and HACTS-DP demonstrated the ability to effectively handle these OOD situations, thereby improving the model’s generalization capabilities.
@@ IV-D Experimental Setup on Reinforcement Learning
To investigate whether HACTS can support human-in-the-loop online learning, we implement HACTS-RLPD, which combines HACTS with RLPD [1], a sample-efficient off-policy reinforcement learning algorithm that uses human demonstrations (expert data) and corrections (HACTS data) during online training. The training proceeds in three stages. In the first stage, similar to HIL-SERL [13], we train a binary classifier based on a pretrained ResNet-10 [6] to process images from the front and wrist cameras. Using HACTS, we collect 200 positive data points indicating task completions and around 1,000 negative data points to train the reward classifier. In the second stage, we gather 20 trajectories of expert data via our system and use them to train a behavior cloning (BC) policy. The BC policy takes concatenated ResNet-10 embeddings, gripper status and proprioceptive states—defined by the relative 6D end-effector pose compared to its initial pose—as input, and the action space corresponds to the end-effector’s 6D movement and the gripper status. In the final stage, we load the pretrained actor network and fine-tune it with RLPD, collecting short, frequent human corrections via HACTS. The reward classifier is queried at every step during online training, returning 1 if the task is completed and 0 otherwise.
As shown in Figure 3, we train and evaluate HACTS-RLPD on a UR5 workstation with our HACTS system, where the environment updates at 10 Hz. As shown in Figure 5, we design a CloseBin task where the robot is required to close a trash bin that is randomly placed and oriented. This task is challenging because precise motion control is required based on the bin cover’s position and orientation.

### v2

@@ I Introduction
However, current teleoperation systems are limited by their hardware, such as VR [khazatsky2024droid], exoskeletons [fang2024airexo], and motion-capture technologies [darvish2023teleoperation], which are designed for unilateral control. In these systems, the human operator can send commands to the robot but receives no real-time feedback, a limitation when the robot is performing autonomous tasks and requires human intervention. Without feedback, the robot may make unintended, sudden movements, disrupting task continuity and performance. This is akin to a driver of an autonomous vehicle attempting to intervene while the vehicle is in motion—without a mechanism to sync the vehicle’s wheel movement with the driver’s input, the intervention could lead to a loss of control [zhang2020toward]. Recent work [kobayashi2025alpha, liu2025factr, toedtheide2023force] has proposed bilateral control systems that integrate tactile and force feedback into teleoperation hardware. However, these systems require the follower device to generate precise force or tactile sensing information—capabilities that are not widely available on most robotic platforms. Thus, there is a clear need for a teleoperation system offering bilateral joint synchronization between the teleoperation system and the robot, ensuring smooth transitions and more reliable interventions, as well as being compatible with a broad range of robots.
To address this limitation, in this paper, we propose HACTS (Human-As-Copilot Teleoperation System), a low-cost but effective solution that enables bilateral, real-time joint synchronization between the robot arm and the teleoperation hardware. The HACTS system is implemented using only 3D-printed components and low-cost, off-the-shelf motors, making it both affordable and scalable. This simple yet effective hardware setup allows for continuous joint feedback from the robot to the teleoperation hardware, enabling more intuitive and controlled intervention when required. It offers a robust, seamless connection between the operator and the robot arm, making it ideal for human-in-the-loop applications.
@@ II-A Teleoperation for robot manipulation learning
Despite recent advances in robot manipulation driven by AI, fully autonomous solutions remain far from achieving socially and physically competent robot behaviors. This is why teleoperated robots, serving as physical avatars for human workers on-site, present a more practical solution [darvish2023teleoperation, tong2024advancements]. Furthermore, with the rise of learning-based, data-driven approaches in robotic control, teleoperation has become an essential system for gathering manipulation data [wu2024robomind, khazatsky2024droid]. Traditional teleoperation systems, such as VR [zhang2018deep], joystick controller [Kinova], Phone-based [mandlekar2018roboturk], and 3D Spacemouse [luo2024precise], can only control the end-effector pose of the robot arm that heavily relies on IK/FK and motion planning, which limits the expression of human operators. More recent approaches have highlighted the potential of low-cost, kinematically equivalent controllers, such as ALOHA [zhao2023learning] and Gello [wu2024gello], which enable more dexterous and precise control, providing valuable data for robot learning. However, these systems have primarily focused on unilateral control, restricting the full potential of the hardware. Several works have attempted to develop bilateral control systems for robotic teleoperation [toedtheide2023force, liu2025factr, kobayashi2025alpha], focusing on providing force or tactile feedback to human operators. However, these complex bilateral systems require expensive follower hardware and carefully designed algorithms to handle force or tactile feedback, which are not accessible on most robotic platforms like the UR5. Although the Bi-ACT [kobayashi2025alpha] system could operate without force feedback, it has only been tested on identical leader-follower setups, and its applicability to larger, more common robotic platforms remains unknown.
In this work, we propose a similar low-cost, kinematically equivalent device that incorporates bilateral control, enabling humans to act as copilots and intervene when needed. We demonstrate that even a simple bilateral position synchronization, which is compatible with almost all robotic systems, can significantly enhance control precision and data collection capabilities. Our approach highlights the potential of more accessible, cost-effective teleoperation systems for improving robot learning and human-robot interaction. Concurrent work RoboCopilot [wu2025robocopilot] introduces a similar human-in-the-loop teleoperation setup, whereas our work further demonstrates a reinforcement learning implementation based on the proposed HACTS framework.
@@ II-B Learning from human demonstration and correction
Learning from Demonstration (LfD) is a key paradigm in robot learning, enabling robots to acquire new skills by learning from human demonstrations rather than through explicit programming. Among the notable advancements in LfD are Vision-Language-Action (VLA) models [ma2024survey, black2024pi_0, wen2025tinyvla, wen2025dexvla]. These approaches facilitate learning multimodal human demonstrations, increasingly leveraging larger model architectures with more parameters. As a result, robots can benefit from richer contextual information, which improves task performance and aligns with the scaling laws observed in large language models. However, the process of collecting human demonstrations can be expensive and time-consuming, with previous work primarily focusing on success trajectories [wu2024robomind, khazatsky2024droid]. In contrast, our HACTS, with its simple bilateral control, allows for human intervention and generates more diverse types of data. Recent research has underscored the importance of failure data in training [wu2025swbt], and we believe our system can significantly contribute to advancing this line of work by supporting the collection of such critical data for more robust learning.
The other group of methods that can benefit from HACTS is Human-In-The-Loop (HITL) learning, which refines robot behavior through interactive feedback and enables policy adjustments based on human corrections. HITL methods can be categorized into passive human intervention [ross2011reduction, luo2023human], where humans assist periodically or as determined by functions, and active human intervention [kelly2019hg, NEURIPS2023_f57ffe47, luo2024precise], where humans decide when to intervene. Recently, approaches such as HACO [NEURIPS2023_f57ffe47] and HIL-SERL [luo2024precise] leverage active human assistance to enhance RL exploration and ensure safety during training, improving sample efficiency and exhibiting robust behaviors compared to classical RL and imitation methods like HG-Dagger [kelly2019hg] and IWR [mandlekar2020human]. However, these approaches require agile hardware platforms for flexible, timely intervention. In contrast to previous teleoperation systems, our HACTS offers more dexterous, precise control during human intervention, which we believe will inspire further innovation in HITL RL, especially for real-world online robot learning.
@@ III-A Overview
We illustrate the key advancements of HACTS in Fig. 2. HACTS operates as a leader-follower system with bilateral position control, enabling both leader-to-follower and follower-to-leader synchronization. In leader-to-follower synchronization, HACTS functions similarly to systems like ALOHA [zhao2023learning] and Gello [wu2024gello], facilitating the rapid and precise collection of offline demonstration data. Conversely, in follower-to-leader synchronization, where model inference drives the robot arm, HACTS enables the human operator to act as a copilot, seamlessly intervening when necessary to provide online action-correction data. Below, we outline key hardware and software design principles in HACTS.
@@ III-B Hardware Design
Low-Cost. One of our primary goals is to demonstrate that a capable teleoperation system can be built at an affordable price, making it accessible for large-scale deployment. To achieve this, we opted for servo motors rather than more expensive robotic modules, which are commonly used in more complex systems [darvish2023teleoperation]. The use of servo motors allows us to implement both passive position reading and active position control at a significantly lower cost. Previous works, including ALOHA [zhao2023learning], Gello [wu2024gello], and FACTR [liu2025factr], have shown the impressive performance and stability of low-cost DYNAMIXEL motors, which further supports our design choice.
@@
Easy-to-Use. Building upon prior work that demonstrated the effectiveness of miniature [wu2024gello], kinematically equivalent devices for teleoperation, we aimed to create hardware that is intuitive and user-friendly, even for non-experts. This simplicity is key to ensuring that the system can be scaled for broader use, i.e., quickly adapt to a new robot. The kinematically equivalent structure is derived from the Denavit–Hartenberg parameters, with the lengths scaled by a factor to maintain the system’s usability while optimizing for lightweight design. Note that these components can easily be extended to other robot arms since we only need to synchronize position information, which is generally available for most robotic arms.
@@ IV Experiments
We evaluate the efficacy of HACTS by addressing four key research questions: Q1: Can human intervention data collected via the HACTS system serve as a viable substitute for an equivalent volume of direct teleoperation data? Q2: Can the action-correction data in HACTS further improve the model’s success rate under in-distribution (ID) scenarios? Q3: Can the action-correction data in HACTS enhance the model’s generalization capability for out-of-distribution (OOD) scenarios? Q4: Can HACTS further support human-in-the-loop online reinforcement learning?
To investigate these questions, we developed real-world robotic arm environments, designed a range of manipulation tasks, and established diverse experimental setups incorporating the two primary learning paradigms: imitation learning (IL) and reinforcement learning (RL).
@@ IV-A Hardware Setup
As shown in Fig. 3, we built HACTS on two workstations, each equipped with one wrist-mounted 336 camera, and one or multiple 336L cameras. As illustrated in Fig. 3, these cameras are positioned at different angles to capture a comprehensive set of visual inputs from varying perspectives. All components—including the cameras, the UR5 robot with Robotiq 2f-85 gripper, the HACTS system, and the policy model—are integrated into a single RTX 4090 workstation, ensuring robust computational performance for real-time data processing. To demonstrate the feasibility and effectiveness of HACTS, we implemented the hardware using the UR5 kinematic model, and reused parts of the structure design in [wu2024gello] but modified them to fit the XL430 motors. The whole cost for HACTS hardware is lower than $\mathdollar 300$, and we listed all components in Tab. I.
@@ IV-B Experimental Setup on Imitation Learning
Task Design. To evaluate the use of the HACTS within the framework of imitation learning, we designed three real-world robotic tasks. These tasks required the robotic arm to learn and execute complex manipulation actions, such as grasping, placing, pressing, flipping, and precisely identifying the location of objects. The objects used in these tasks are shown in Fig. 4. OpenBox: Accurately locate a switch of a box and press it to open the box. SteamBun: Grasp a steamed bun from a plate and place it into a steamer. UprightMug: Flip an overturned mug upright, then place it back in a designated central position.
Algorithm Instances. We focus on popular visuomotor policy learning techniques, where the robot learns to perform tasks based on visual observations and proprioception. The input consists of the RGB image data and the robot proprioceptive states, while the output is the end-to-end robotic control signals, such as joint positions. Imitation learning methods allow the robot to learn directly from human demonstrations, mimicking the demonstrated actions to perform similar tasks autonomously. We employed two state-of-the-art imitation learning methods for our experiments:
  • Action Chunking Transformer (ACT) [zhao2023learning]: ACT predicts a sequence of actions and leverages temporal ensemble technology to generate smooth and precise actions.
  • Diffusion Policy (DP) [chi2023diffusion]: DP uses a diffusion model, which is expressive and can model multi-modal action patterns, to improve the success rates.
Evaluation Protocol. For each task, we initially collected 100 expert trajectories through teleoperation. From these, we randomly chose 50 expert trajectories to train the preliminary ACT and DP models, referred to as “pre-ACT//DP”. Additionally, we trained the ACT and DP models using the full set of 100 expert trajectories as the baseline models, referred to as “full-ACT//DP”. Further, to answer the Q1, Q2 and Q3, we designed four different experimental settings. For each various experimental setting, we used HACTS to collect 50 human intervention trajectories (denoted as HACTS trajectories) with using the ”pre-ACT//DP” to execute tasks. Then we mixed these 50 HACTS trajectories with the initial 50 random expert trajectories to train the refined ACT and DP models, denoted as “HACTS-ACT//DP”. The four experimental settings are as follows:
@@ IV-C Experimental Results on Imitation Learning
Tab. II presents the results of ACT on the three experimental settings including EADC, FCID, and ODSS for all three tasks, evaluated in terms of success rate. Tab. III presents the results of the DP algorithm on the setting ODDS.
Answer for Question 1. We focused on the comparison between HACTS-ACT and full-ACT in the EADC setting in Tab. II. Although trained on the same amount of data, HACTS-ACT achieved higher success rates than full-ACT in all three tasks, due to including more demonstrations in scenarios where pre-ACT models often failed, such as picking actions. This indicates that the HACTS data can effectively replace expert teleoperation data without any performance drop.
Answer for Question 2. We examined whether the performance of HACTS-ACT improves in the FCID setting. The results presented in Tab. II reveal that HACTS-ACT not only significantly enhanced the success rates of pre-ACT across all three tasks, but also consistently outperformed full-ACT. For instance, in the UprightMug task, HACTS-ACT achieved a remarkable success rate of 70%, compared to just 10% for pre-ACT and 40% for full-ACT. This improvement can be attributed to the ability of HACTS-ACT to specifically gather corrective data targeted at failed cases, whereas full-ACT merely increases the volume of data.
Answer for Question 3. We analyzed the performance of HACTS-ACT under the ODSS setting in Tab. II and the ODDS setting in Tab. III. Notably, we found that the success rates of pre-ACT, pre-DP, full-ACT, and full-DP are all 0% in these scenarios, indicating that they are unable to handle situations that are unseen in the training dataset. However, upon incorporating HACTS-generated action-correction data, both HACTS-ACT and HACTS-DP demonstrated the ability to effectively handle these OOD situations, thereby improving the model’s generalization capabilities.
@@ IV-D Experimental Setup on Reinforcement Learning
To investigate whether HACTS can support human-in-the-loop online learning, we implement RLPD-HACTS, which combines HACTS with RLPD [ball2023efficient], a sample-efficient off-policy reinforcement learning algorithm that uses human demonstrations (expert data) and corrections (HACTS data) during online training. The training proceeds in three stages. In the first stage, similar to HIL-SERL [luo2024precise], we train a binary classifier based on a pretrained ResNet-10 [he2016deep] to process images from the front and wrist cameras. Using HACTS, we collect 200 positive data points indicating task completions and around 1,000 negative data points to train the reward classifier. In the second stage, we gather 20 trajectories of expert data via our system and use them to train a behavior cloning (BC) policy. The BC policy takes concatenated ResNet-10 embeddings, gripper states and proprioceptive states—defined by the relative 6D end-effector pose compared to its initial pose—as input, and the action space corresponds to the end-effector’s 6D movement and the gripper states. In the final stage, we load the pretrained actor network and fine-tune it with RLPD, collecting short, frequent human corrections via HACTS. The reward classifier is queried at every step during online training, returning 1 if the task is completed and 0 otherwise.
As shown in Fig. 3, we train and evaluate RLPD-HACTS on a UR5 workstation with our HACTS system, where the environment updates at 10 Hz. As shown in Fig. 5, we design a CloseBin task where the robot is required to close a trash bin that is randomly placed and oriented. This task is challenging because precise motion control is required based on the bin cover’s position and orientation.